Write down all elements of quotient group Z18/<6>. Is any element of order 5?
1
Expert's answer
2020-10-13T18:31:48-0400

Obviously

"^{\mathbb{Z}_{18}}\!/\!_{<6>}= \mathbb{Z}_6=\{0,1,2,3,4,5\}"

There are no elements of order 5 here because "ord(\mathbb{Z}_6)=6" and order of any element must be a divisor of the group order, but "5\nmid6"